Math Club is having a meeting. If student activities has and event at the same time, I estimate that 12 students will come to Math Club. If student activities does not have an event at the same time, I estimate 30 students will come. The probability that student activities will have an event at the same time is 0.35. Find the expected value.
Answer #1

Expected value = [P(student activity occurring) * (No. of students coming given there is a student activity)] + [P(student activity not occurring) * (No. of students coming given there is no student activity)]

Expected Value = (0.35 * 12) + (1 - 0.35) * 30 = 4.2 + 19.5 = 23.7